Name:

SS#:

Test 1

Instructions: Answer all seven problems, each one on a separate answer sheet. Put the number of the problem you are doing at the top of the answer sheet.

All problems are worth the same amount.

1. A group of students are polled regarding the number of tacos they consumed last month. Their responses were: 4, 15, 3, 0, 6, 9, 5.

(a) Find the sample mean.

(b) Find the sample variance.

(c) Find the sample standard deviation.

2. Martin is playing Dungeons and Dragons with some of his friends. He is using two dice, both of which are fair, but one has 7 sides and the other has 4. He must roll a number greater than 9 to defeat a troll. What is the probability that he suceeds?

Bonus: What is the probability that Martin has no life?

3. Fill in all 8 regions in the Venn diagram, given the following information:

P (ABC) = .1 P (BC) = .2 P (\overline{A}C) = .3 P (AC) = .15 P (AB) = .2 P (A ∪ C) = .6 P (B) = .6